# MoColl: Agent-Based Specific and General Model Collaboration for Image Captioning

### v1

@@ Abstract
Image captioning is a critical task at the intersection of computer vision and natural language processing, with wide-ranging applications across various domains. For complex tasks such as diagnostic report generation, deep learning models require not only domain-specific image-caption datasets but also the incorporation of relevant general knowledge to provide contextual accuracy. Existing approaches exhibit inherent limitations: specialized models excel in capturing domain-specific details but lack generalization, while vision-language models (VLMs) built on large language models (LLMs) leverage general knowledge but struggle with domain-specific adaptation. To address these limitations, this paper proposes a novel agent-enhanced model collaboration framework, which we called MoColl, designed to effectively integrate domain-specific and general knowledge. Specifically, our approach is to decompose complex image captioning tasks into a series of interconnected question-answer subtasks. A trainable visual question answering (VQA) model is employed as a specialized tool to focus on domain-specific visual analysis, answering task-specific questions based on image content. Concurrently, an LLM-based agent with general knowledge formulates these questions and synthesizes the resulting question-answer pairs into coherent captions. Beyond its role in leveraging the VQA model, the agent further guides its training to enhance its domain-specific capabilities. Experimental results on radiology report generation validate the effectiveness of the proposed framework, demonstrating significant improvements in the quality of generated reports.
@@ 1 Introduction
On the other hand, general models utilizing Vision-Language Models (VLMs) (OpenAi, 2023; Liu et al., 2023c; Wang et al., 2023; Bai et al., 2023b; Lu et al., 2024) present a compelling alternative (see fig. 1(b) for an illustration). VLMs, built upon the foundation of Large Language Models (LLMs), are extensively versatile, highly scalable, and imbued with substantial background knowledge. Therefore, these models have the ability to generate informative captions across a wide range of domains following instructions. However, general models face significant challenges when applied to specific-domain tasks. Two common strategies are used to adapt general models: In-Context Learning (ICL) and fine-tuning. ICL improves domain knowledge but does not enhance the model’s ability to perform domain-specific tasks. On the other hand, fine-tuning can lead to catastrophic forgetting, where domain-specific training may degrade the model’s ability to retain general knowledge (McCloskey and Cohen, 1989; Luo et al., 2023). As a result, neither approach effectively combines general knowledge with domain expertise for complex captioning tasks.
@@ 2.2 Vision-Language Model
Vision-Language Models (VLMs) such as GPT-4V (OpenAi, 2023), LLaVA (Liu et al., 2023c, b), DeepSeek-VL (Lu et al., 2024), CogVLM (Wang et al., 2023), and Qwen-VL (Bai et al., 2023b), are usually based on LLMs such as GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023), DeepSeek LLM (Bi et al., 2024), Vicuna (Chiang et al., 2023), and Qwen-LM (Bai et al., 2023a) respectively. These models integrate visual and textual data, treating tokenized images as another form of language input. This allows them to leverage background knowledge, which in turn provides comprehensive understanding and generation capabilities. However, despite their extensive background knowledge, general VLMs often lack the granularity and specificity in domain expertise that specialized models possess, which manifests as a less detailed grasp of specialized knowledge.
@@ 2.3 Large Language Model as Agent
Recent research, such as MetaGPT (Hong et al., 2023) and AutoGPT (Significant Gravitas, ), has found that LLMs can be employed as agents, acting as core controllers for executing complex tasks. In particular, their application extends across multiple domains of functionality, each illustrating the versatility of their capabilities as autonomous systems.
@@
One of the most significant advancements in utilizing LLMs as agents is their role as optimizers. LLMs excel in optimizing prompts (Yang et al., 2024), where they generate and refine input queries to maximize the effectiveness and relevance of their responses. What’s more, LLMs have also demonstrated a remarkable ability in refining machine learning models by tuning hyperparameters (Zhang et al., 2023; Liu et al., 2024), suggesting architectural improvements (Zheng et al., 2023), improving loss functions (Song et al., 2023; Wu et al., 2023), and even guiding exploration in reinforcement learning (Du et al., 2023). Building on this work, we propose to innovate further by leveraging LLMs to optimize data, i.e., dynamically generating and refining training datasets, particularly through the creation of targeted visual-question-answer data that address specific weaknesses of VQA models.
@@ 3.1 Model Collaboration Framework
$c_{i}^{*}=\text{Agent}_{\text{caption}}(\{(q_{i},a_{i})\}_{i=1}^{m^{*}}).$ (3)
Thus, we get the generated caption $c_{i}^{*}$.
@@
1:the image $x$, the agent $\text{Agent}_{\text{question}}$ and $\text{Agent}_{\text{caption}}$, and the VQA model
@@
3:     $(q_{j},\ \text{stop\_sign})\leftarrow\text{Agent}_{\text{question}}(\{q_{k},a_{k}\}_{k<j})$triangleright questioning
4:     if stop_sign is true then
5:         break
@@
7:     $a_{j}\leftarrow\text{VQA}(x,q_{j})$triangleright answering
@@
9:$c^{*}\leftarrow\text{Agent}_{\text{caption}}(\{q_{k},a_{k}\}_{k<j})$triangleright captioning
10:return $\{\{q_{k},a_{k}\}_{j=1}^{m_{i}},c^{*}\}$
@@ 3.2 Training Procedure
The training procedure consists of the following 2 stages. The first stage is the warm-up stage, where we align the image features to their textual word embeddings in the pre-trained LM. The second stage is the main training stage called agent-guided tuning, where we fine-tune the VQA model with visual-question-answer data generated by interaction between the agent and VQA model and selected by the agent. It is detailed as follows and illustrated in fig. 2.
@@ 3.2.1 Stage 1: Warm up
3.2.2 Stage 2: Agent-Guided Tuning.
@@
This step aims to synthesize many synthetic question-answer pairs from image-caption pairs in the training dataset. Specifically, we simulate the questioning step of the model collaboration framework (as detailed in section 3.1) in which the agent continuously asks questions to the VQA model and receives answers (see eq. 1 and eq. 2). Similarly, the iteration will not stop until the agent ensures that enough image information has been collected through these question-answer pairs or achieves a stopping criterion. The inference process is in lines 1-7 of algorithm 1.
@@
This step aims to select high-quality synthetic question-answer pairs and thus build a VQA dataset. Inspired by recent work on selecting synthetic data, we propose that the agent can be used to improve the performance of the VQA model via selected synthetic data. Specifically, we instruct the agent to select correct question-answer pairs according ground truth captions. Similar to the process of a teacher looking at a standard answer to correct a student’s error, the selection agent can find out inappropriate questions from the above questioning agent and inaccurate answers from the VQA model. Accordingly, the selection agent designs new questioning prompts and generates additional question-answer pairs
$\{(q_{i},a_{i})\mid\text{Agent}_{\text{select}}(q_{ij},a_{ij},c_{i})=\text{True}\}$ (4)
@@
This step aims to fine-tune the VQA model with the high-quality VQA dataset. It improves the performance of the VQA model and ensures that the tool remains effective and adaptable. Specifically, we use auto-regressive loss, i.e., $p(a_{j}\mid x_{j},q_{j};\text{VQA})$, where $(x_{j},q_{j},a_{j})$ is the selected image-question-answer tuple. The visual encoder weights are still frozen, and the LM and the projection network can be updated.
We summarize the agent-guided tuning process in algorithm 2 and illustrate it in fig. 2(c).
@@
1:an image captioning dataset $\mathcal{D}=\{(x_{i},c_{i})\}_{i=1}^{N}$, an initialized $\text{VQA}_{\theta}$ model (an aligned VLM), the agent $\text{Agent}_{\text{question}}$ and $\text{Agent}_{\text{select}}$
2:// Step 1: Inference (see algorithm 1). $\triangleright$ MoColl framework without captioning
3:Initialize $\text{Memories}\leftarrow\text{Set}()$
4:for $i\leftarrow 1\ \text{to}\ N$ do
5:     Sample $(x_{i},c_{i})\sim\mathcal{D}$
@@
7:         $(q_{ij},\text{stop\_sign})\leftarrow\text{Agent}_{\text{question}}(\{(q_{ik},a_{ik})\}_{k<j})$triangleright questioning
8:         if stop_sign is true then
9:              break
10:         end if
11:         $a_{ij}\leftarrow\text{VQA}(q_{ij},x_{i})$triangleright answering
12:         Memories.add($(\{x_{i},q_{ij},a_{ij},c_{i}\})$
13:     end for
14:end for
15:// Step 2: Select according to the ground truth caption.
[…]
18:     if $\text{Agent}_{\text{select}}(q_{ij},a_{ij},c_{i})$ is true then $\triangleright$ selecting
[…]
20:     end if
21:end for
22:// Step 3: Train the VQA model.
[…]
25:     $\mathcal{L}(\theta)=\frac{1}{B}\sum_{j=1}^{B}\log p(a_{j}\mid x_{j},q_{j};\text{VQA}_{\theta})$triangleright auto-regressive loss
26:     update $\theta$ through an optimizer, e.g., ADMM
27:end for
28:return $\text{VQA}_{\theta}$ model
@@ 4.1 Implementation
We detail the specialized VQA tool and LLM-based agent employed in our framework. For the VQA model, we use LLaVA (Liu et al., 2023c), a visual-language conversation model, as our VQA model. The architecture remains unchanged, consisting of a frozen image encoder, a projection network, and a pre-trained causal Language Model (LM). Specifically, we use CLIP Vision Encoder (Radford et al., 2021) as the image encoder, which outputs a vector representing the features of the input image. The image features are then mapped into visual tokens via a randomly initialized three-layer Multi-Layer Perceptron (MLP), which serves as an intermediary, projecting the image information into a "foreign language" understood by the LM. The pre-trained LM is Llama-3.1-8B (Dubey et al., 2024). For the agent, we adopt DeepSeek-V2.5 (DeepSeek-AI, 2024). Furthermore, we demonstrate our model collaboration (MoColl) framework with two tools: the aligned VLM and the tuned VQA model. The former is trained only through a warm-up stage, while the latter is trained from the aligned VLM through an agent-guided tuning stage. Additional implementation details can be found in section B.3.
@@
We evaluate our method against three categories of baseline approaches, as depicted in fig. 1. (1) Specialized encoder-decoder models: R2Gen (Chen et al., 2020), Joint-TriNet (Yang et al., 2021), R2GenCMN (Chen et al., 2021), XProNet (Wang et al., 2022a), and M2KT (Yang et al., 2023). (2) General vision language models: LLaVA-Med (Li et al., 2023), and LLaVA-1.5-7b (Liu et al., 2023b), DeepSeek-VL2 (Lu et al., 2024), Qwen2-VL-72B-Instruct (Bai et al., 2023b) (denoted as Qwen2-VL), our aligned VLM, and supervised fine-tuned VLM. (3) Model collaboration framework with VQA models: LLaVA-Med and LLaVA-1.5-7b. Implementation details of each baseline method are provided in the section B.4.
@@ A.1 Radiology Report Generation for Chest X-ray
To incorporate medical knowledge, some studies integrated clinical ontologies and knowledge bases into their models. For example, Li et al. (2019) introduced a model that aligns image features with medical concepts, ensuring the use of correct terminology and facilitating the generation of more informative reports. Additionally, Yang et al. (2023) adopted a multi-task knowledge transfer mechanism to enhance the generation process by simultaneously learning disease classification and report generation tasks.
Despite these advancements, existing models often struggle to fully leverage both the general language understanding capabilities of large-scale models and the domain-specific expertise required for medical reporting. Challenges such as generating coherent narratives, accurately reflecting clinical findings, and adapting to varied patient cases remain. Our method is capable of addressing these challenges by combining the expansive knowledge and language generation abilities of Large Language Models (LLMs) with the specialized visual processing capabilities of a domain-specific VQA model. Therefore, we use the radiology report generation task as an example to run experiments, demonstrating the effectiveness of our approach.

### v3

@@ Abstract
Image captioning is a critical task at the intersection of computer vision and natural language processing, with wide-ranging applications across various domains. For complex tasks such as diagnostic report generation, deep learning models require not only domain-specific image-caption datasets but also the incorporation of relevant general knowledge to provide contextual accuracy. Existing approaches exhibit inherent limitations: specialized models excel in capturing domain-specific details but lack generalization, while vision-language models (VLMs) built on large language models (LLMs) leverage general knowledge but struggle with domain-specific adaptation. To address these limitations, this paper proposes a novel agent-enhanced model collaboration framework, which we call MoColl, designed to effectively integrate domain-specific and general knowledge. Specifically, our approach is to decompose complex image captioning tasks into a series of interconnected question-answer subtasks. A trainable visual question answering (VQA) model is employed as a specialized tool to focus on domain-specific visual analysis, answering task-specific questions based on image content. Concurrently, an LLM-based agent with general knowledge formulates these questions and synthesizes the resulting question-answer pairs into coherent captions. Beyond its role in leveraging the VQA model, the agent further guides its training to enhance its domain-specific capabilities. Experimental results on radiology report generation validate the effectiveness of the proposed framework, demonstrating significant improvements in the quality of generated reports.
@@ 1 Introduction
On the other hand, general models utilizing Vision-Language Models (VLMs) (OpenAi, 2023; Liu et al., 2023b; Wang et al., 2023; Bai et al., 2023b; Lu et al., 2024) present a compelling alternative (see fig. 1(b) for an illustration). VLMs, built upon the foundation of Large Language Models (LLMs), are extensively versatile, highly scalable, and imbued with substantial background knowledge. Therefore, these models have the ability to generate informative captions across a wide range of domains following instructions. However, general models face significant challenges when applied to specific-domain tasks. Two common strategies are used to adapt general models: In-Context Learning (ICL) and fine-tuning. ICL improves domain knowledge but does not enhance the model’s ability to perform domain-specific tasks. On the other hand, fine-tuning can lead to catastrophic forgetting, where domain-specific training may degrade the model’s ability to retain general knowledge (McCloskey and Cohen, 1989; Luo et al., 2023). As a result, neither approach effectively combines general knowledge with domain expertise for complex captioning tasks.
@@ 2.2 Vision-Language Model
Vision-Language Models (VLMs) such as GPT-4V (OpenAi, 2023), LLaVA (Liu et al., 2023b, 2024b), DeepSeek-VL (Lu et al., 2024), CogVLM (Wang et al., 2023), and Qwen-VL (Bai et al., 2023b), are usually based on LLMs such as GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023), DeepSeek LLM (Bi et al., 2024), Vicuna (Chiang et al., 2023), and Qwen-LM (Bai et al., 2023a) respectively. These models integrate visual and textual data, treating tokenized images as another form of language input. This allows them to leverage background knowledge, which in turn provides comprehensive understanding and generation capabilities. However, despite their extensive background knowledge, general VLMs often lack the granularity and specificity in domain expertise that specialized models possess, which manifests as a less detailed grasp of specialized knowledge.
@@ 2.3 Large Language Model as Agent
Recent research, such as MetaGPT (Hong et al., 2024) and AutoGPT (Significant Gravitas, ), has found that LLMs can be employed as agents, acting as core controllers for executing complex tasks. In particular, their application extends across multiple domains of functionality, each illustrating the versatility of their capabilities as autonomous systems.
@@
One of the most significant advancements in utilizing LLMs as agents is their role as optimizers. LLMs excel in optimizing prompts (Yang et al., 2024), where they generate and refine input queries to maximize the effectiveness and relevance of their responses. What’s more, LLMs have also demonstrated a remarkable ability in refining machine learning models by tuning hyperparameters (Zhang et al., 2023; Liu et al., 2024c), suggesting architectural improvements (Zheng et al., 2023), improving loss functions (Song et al., 2023; Wu et al., 2023), and even guiding exploration in reinforcement learning (Du et al., 2023). Building on this work, we propose to innovate further by leveraging LLMs to optimize data, i.e., dynamically generating and refining training datasets, particularly through the creation of targeted visual-question-answer data that address specific weaknesses of VQA models.
@@ 3.1 Model Collaboration Framework
$\hat{c}_{i}=\text{Agent}_{\text{caption}}(\{(q_{ij},a_{ij})\}_{j}).$
Thus, we get the generated caption $\hat{c}_{i}$.
@@
1:the image $x$, the agent $\text{Agent}_{\text{question}}$ and $\text{Agent}_{\text{caption}}$, the VQA model, the maximum number of questions $M$
@@
3:     $q_{j}\leftarrow\text{Agent}_{\text{question}}(\{(q_{k},a_{k})\}_{k<j})$triangleright question
4:     if $q_{j}$ is None then
5:         break $\triangleright$ enough information
@@
7:     $a_{j}\leftarrow\text{VQA}(x,q_{j})$triangleright answer
@@
9:$\hat{c}\leftarrow\text{Agent}_{\text{caption}}(\{(q_{k},a_{k})\}_{k<j})$triangleright caption
10:return $\{\{(q_{k},a_{k})\}_{k<j},\hat{c}\}$
@@ 3.2 Training Procedure
The training procedure consists of the following 2 stages. The first stage is the warm-up stage, where we align the image features to their textual word embeddings in the pre-trained LM. The second stage is the main training stage called agent-guided tuning, where we fine-tune the VQA model with visual-question-answer data generated by interaction between the agent and VQA model and selected by the agent. It is detailed as follows and the whole training procedure is illustrated in fig. 2(a).
@@ 3.2.1 Stage 1: Warm up
3.2.2 Stage 2: Agent-Guided Tuning
@@
We summarize the agent-guided tuning process in algorithm 2 and illustrate it in fig. 2(c). Details of these three steps are shown as follows.
@@
This step aims to synthesize many synthetic question-answer pairs from image-caption pairs in the training dataset. Specifically, we simulate the questioning step of the model collaboration framework (as detailed in section 3.1) in which the agent continuously asks questions to the VQA model and receives answers (see eq. 1 and eq. 2). Similarly, the iteration will not stop until the agent ensures that enough image information has been collected through these question-answer pairs or achieves a stopping criterion. The inference process is in lines 1-12 of algorithm 2.
@@
This step aims to select high-quality synthetic question-answer pairs and thus build a VQA dataset. Inspired by recent work on selecting synthetic data, we propose that the agent can be used to improve the performance of the VQA model via selected synthetic data. Specifically, we instruct the agent to select correct question-answer pairs according ground truth captions. Similar to the process of a teacher looking at a standard answer to correct a student’s error, the selection agent can find out inappropriate questions from the above questioning agent and inaccurate answers from the VQA model, and then discards them. Subsequently, it remains high-quality question-answer pairs for the next training step, formalized as
$\{(q_{ij},a_{ij})\mid\text{Agent}_{\text{select}}(q_{ij},a_{ij},c_{i})=\text{True}\}.$
The selection process is in lines 13-18 of algorithm 2.
@@
This step aims to fine-tune the VQA model with the high-quality VQA dataset. It improves the performance of the VQA model and ensures that the tool remains effective and adaptable. Specifically, we use auto-regressive loss, i.e., $p(a_{j}\mid x_{j},q_{j};\text{VQA})$, where $(x_{j},q_{j},a_{j})$ is the selected image-question-answer tuple. The visual encoder weights are still frozen, and the LM and the projection network can be updated. The training process is in lines 19-24 of algorithm 2.
Overall, with the whole training procedure, we not only integrated domain-specific knowledge but also significantly enhanced the VQA model’s capabilities. In the following experimental section, We will conduct a series of experiments to demonstrate the advantages of our methods.
@@
1:an image captioning dataset $\mathcal{D}=\{(x_{i},c_{i})\}_{i=1}^{N}$, an aligned VLM, the agent $\text{Agent}_{\text{question}}$ and $\text{Agent}_{\text{select}}$, the maximum number of questions $M$ // Step 1: Infer (see algorithm 1) without captioning.
2:Initialize Memories $\mathcal{M}\leftarrow\text{Set}()$
3:for $i\leftarrow 1\ \text{to}\ N$ do
4:     Sample $(x_{i},c_{i})\sim\mathcal{D}$
5:     for $j\leftarrow 1\ \text{to}\ M$ do
6:         $q_{ij}\leftarrow\text{Agent}_{\text{question}}(\{(q_{ik},a_{ik})\}_{k<j})$triangleright question
7:         if $q_{ij}$ is None then
8:              break $\triangleright$ enough information
9:         end if
10:         $a_{ij}\leftarrow\text{VLM}(q_{ij},x_{i})$triangleright answer
11:         Add $(x_{i},q_{ij},a_{ij},c_{i})$ into $\mathcal{M}$
12:     end for
13:end for// Step 2: Select according to the ground truth caption.
14:Initialize the VQA dataset $\mathcal{D}^{\text{VQA}}\leftarrow\text{Set}()$
15:for $(x_{i},q_{ij},a_{ij},c_{i})\ \text{in}\ \mathcal{M}$ do
16:     if $\text{Agent}_{\text{select}}(q_{ij},a_{ij},c_{i})$ is true then
17:         Add $(x_{i},q_{ij},a_{ij})$ into $\mathcal{D}^{\text{VQA}}$triangleright select
18:     end if
19:end for// Step 3: Train the VQA model.
20:Initialize the VQA model $\text{VQA}_{\theta}\leftarrow\text{VLM}$
@@
22:     Sample a mini-batch $\{(x_{j},q_{j},a_{j})\}_{j=1}^{B}\sim\mathcal{D}^{\text{VQA}}$
23:     $\mathcal{L}(\theta)=\frac{1}{B}\sum_{j=1}^{B}\log p(a_{j}\mid x_{j},q_{j};\text{VQA}_{\theta})$triangleright auto-regressive loss
24:     Update $\theta$ through $\mathcal{L}(\theta)$ with an optimizer, e.g., ADMM
25:end for
26:return $\text{VQA}_{\theta}$ model
@@ 4.1 Implementation
We detail the specialized VQA tool and LLM-based agent employed in our framework. For the VQA model, we use LLaVA (Liu et al., 2023b), a visual-language conversation model, as our VQA model. The architecture remains unchanged, consisting of a frozen image encoder, a projection network, and a pre-trained causal Language Model (LM). Specifically, we use CLIP Vision Encoder (Radford et al., 2021) as the image encoder, which outputs a vector representing the features of the input image. The image features are then mapped into visual tokens via a randomly initialized three-layer Multi-Layer Perceptron (MLP), which serves as an intermediary, projecting the image information into a "foreign language" understood by the LM. The pre-trained LM is Llama-3.1-8B (Dubey et al., 2024). For the agent, we adopt DeepSeek-V2.5 (Liu et al., 2024a). Furthermore, we demonstrate our model collaboration (MoColl) framework with two tools: the aligned VLM and the tuned VQA model. The former is trained only through a warm-up stage, while the latter is trained from the aligned VLM through an agent-guided tuning stage. Additional implementation details can be found in section B.3.
@@
We evaluate our method against three categories of baseline approaches, as depicted in fig. 1. (1) Specialized encoder-decoder models: R2Gen (Chen et al., 2020), Joint-TriNet (Yang et al., 2021), R2GenCMN (Chen et al., 2021), XProNet (Wang et al., 2022a), and M2KT (Yang et al., 2023). (2) General vision language models: LLaVA-Med (Li et al., 2023), and LLaVA-1.5-7b (Liu et al., 2024b), DeepSeek-VL2 (Lu et al., 2024), Qwen2-VL-72B-Instruct (Bai et al., 2023b) (denoted as Qwen2-VL), our aligned VLM, and supervised fine-tuned VLM. (3) Model collaboration framework with VQA models: LLaVA-Med and LLaVA-1.5-7b. Implementation details of each baseline method are provided in the section B.4.
@@ A.1 Radiology Report Generation for Chest X-ray
To incorporate medical knowledge, some studies integrated clinical knowledge bases into their models. For example, Li et al. (2019) introduced a model that aligns image features with medical concepts, ensuring the use of correct terminology and facilitating the generation of more informative reports. Additionally, Yang et al. (2023) adopted a multi-task knowledge transfer mechanism to enhance the generation process by simultaneously learning disease classification and report generation tasks.
Despite these advancements, existing models often struggle to fully leverage both the general knowledge of large-scale models and the domain-specific expertise required for medical reporting. Challenges such as generating coherent narratives, accurately reflecting clinical findings, and adapting to varied patient cases remain. Our method is capable of addressing these challenges by combining the expansive knowledge and language generation abilities of LLMs with the specialized visual processing capabilities of a domain-specific VQA model. Therefore, we use the radiology report generation task as an example to run experiments, demonstrating the effectiveness of our approach.